# Optimal Multi-Distribution Learning

### v2

@@ D.2 Restatement and Proof of Lemma 12
$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$
$=\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
=$\overset{\mathrm{(i)}}{=}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]\right]$
$\overset{\mathrm{(ii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
$=\left(\sum_{i=1}^{k}(n_{i}+m_{i})\right)\widetilde{\mathsf{Rad}}_{\{n_{i}+m_{i}\}_{i=1}^{k}}$
$\overset{\mathrm{(iii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
$\qquad\qquad\qquad\qquad+\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=n_{i}+1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
$=\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$
Here, (i) is valid due to the zero-mean property of $\{\sigma_{i}^{j}\}$, (ii) comes from Jensen’s inequality, and (iii) holds since $\max_{x}(f_{1}(x)+f_{2}(x))\leq\max_{x}f_{1}(x)+\max_{x}f_{2}(x)$. ∎
D.3 Restatement and Proof of Lemma 13
Lemma 13.[Restatement] Consider any $\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for each $i\in[k]$. By taking $w\in\Delta^{k}$ with $w_{i}=\frac{n_{i}}{\sum_{l=1}^{k}n_{l}}$, one has
Proof.
Set $\widetilde{n}=\sum_{i=1}^{k}n_{i}$. Let $\{X_{j}\}_{j=1}^{n}$ be $n$ i.i.d. multinomial random variables with parameter $\{w_{i}\}_{i=1}^{k}$, and take $\widehat{n}_{i}=\sum_{j=1}^{n}\mathds{1}\{X_{j}=i\}$ for each $i\in[k]$. From (44) and Definition 9, it is easily seen that
@@ D.3 Restatement and Proof of Lemma 13
$\mathsf{Rad}_{\widetilde{n}}\big{(}D(w)\big{)}$ $=\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{\widetilde{n}}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{\widetilde{n}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right]\right],$
@@
with probability exceeding $1-\frac{1}{2k}$. By defining $\mathcal{E}$ to be the event where $\widehat{n}_{i}\geq n_{i}/6$ for all $i\in[k]$, we can immediately see from the union bound that
@@
$\mathsf{Rad}_{\widetilde{n}}\big{(}\mathcal{D}(w)\big{)}$ $\geq\mathbb{P}(\mathcal{E})\cdot\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{\widetilde{n}}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{\widetilde{n}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\mid\mathcal{E}\right]\right]\right]$
$\geq\frac{1}{2}\cdot\frac{\sum_{i=1}^{k}\widetilde{n}_{i}}{\widetilde{n}}\widetilde{\mathsf{Rad}}_{\{\widetilde{n}_{i}\}_{i=1}^{k}}$
@@
where the last two inequalities both follow from Lemma 39. This concludes the proof. ∎
D.4 Necessity of Assumption 10
In this section, we will discuss the necessity of Assumption 10 in comparison to the following weaker assumption, where we only assume that the Rademacher complexity on each $\mathcal{D}_{i}$ is well bounded.
@@ D.4 Necessity of Assumption 10
For each $n\geq 1$, there exists a universal constant $C_{n}\geq 0$ such that
$\mathsf{Rad}_{n}(\mathcal{D}_{i})\leq C_{n}$ (110)
for all $i\in[k]$.
@@
Let $w^{0}=[1/k,1/k,\ldots,1/k]^{\top}$. There exists a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that
$\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))\geq\Omega\left(\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})\right)$ (111)
@@
Without loss of generality, we set $\mathcal{Y}=\{0\}$ and $\ell(h,(x,y))=h(x)-y=h(x)$. We can then regard $\mathcal{D}_{i}$ as a distribution over $\mathcal{X}_{i}$ because there is only one element in $\mathcal{Y}$.
We pick $k$ subsets of $\mathcal{X}$ as $\{\mathcal{X}_{i}\}_{i=1}^{k}$. For each $i\in[k]$, we choose the distribution $\mathcal{D}_{i}$ to be an arbitrary distribution supported by $\mathcal{X}_{i}$. On the other hand, we also define $\mathcal{H}_{i}$ to be a set of hypothesis $h(x)=0$ for all $x\notin\mathcal{X}_{i}$ for each $i\in[k]$. For $\{h_{i}\}_{i=1}^{k}$ such that $h_{i}\in\mathcal{H}_{i}$, we define $\mathsf{joint}(\{h_{i}\}_{i=1}^{k})$ to be the hypothesis $h$ such that $h(x)=h_{i}(x),\forall x\in\mathcal{X}_{i},i\in[k]$ and $h(x)=0$ for $x\notin\cup_{i}\mathcal{X}_{i}$. Now we construct the hypothesis set $\mathcal{H}$ as
=$\mathcal{H}=\{\mathsf{joint}(\{h_{i}\}_{i=1}^{k})|h_{i}\in\mathcal{H}_{i},\forall i\in[k]\}.$ (112)
Recall the definition of $\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\{\mathcal{D}_{i}\}_{i=1}^{k})$. Then we have that
$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\{\mathcal{D}_{i}\}_{i=1}^{k})$ $=\mathop{\mathbb{E}}\limits_{\{x_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h(x_{i}^{j})\right]\right]$
$=\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}\max_{h_{i}\in\mathcal{H}_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h_{i}(x_{i}^{j})\right]\right]$ (113)
@@
$=\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}n_{i}\mathsf{Rad}_{n_{i}}(\mathcal{D}_{i}).$
Here (113) is by the definition of $\mathcal{H}$.
By taking $n_{i}=\frac{n}{k}$ for all $i\in[k]$ and Lemma 13, we have
$\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})=\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\{\mathcal{D}_{i}\}_{i=1}^{k})\leq 72\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))$ (114)
By Lemma 32, if we set $C_{n}\sim\sqrt{\frac{d}{n}}$ in Assumption 31, the best upper bound for $\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))\sim\sqrt{\frac{dk}{n}}$, which implies that more samples are needed to learn the mixture distribution $w^{0}$.
Moreover, under the construction in Lemma 32, by further assuming $\min_{h_{i}\in\mathcal{H}_{i}}\mathbb{E}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]=\frac{1}{2}$ for all $i\in[k]$, to find $h$ such that
$\max_{i}\mathbb{E}_{x\sim\mathcal{D}_{i}}[h(x)]\leq\frac{1}{2}+\epsilon,$ (115)
we need to find $h_{i}\in\mathcal{H}_{i}$ such that
$\mathbb{E}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]\leq\frac{1}{2}+\epsilon$ (116)
for all $i\in[k]$.
Following this intuition, we can construct a counter example under Assumption 31 as below.
@@
There exists a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that Assumption 31 holds with $C_{n}=O\left(\sqrt{\frac{d}{n}}\right)$, and it costs at least $\tilde{\Omega}\left(\frac{dk}{\epsilon^{2}}\right)$ samples to find some $h\in\mathcal{H}$ such that
$\max_{i\in[k]}L(h,\mathbf{e}_{i})\leq\min_{h^{\prime}\in\mathcal{H}}\max_{i\in[k]}L(h^{\prime},\mathbf{e}_{i})+\epsilon.$
@@
With the construction in Lemma 32, it suffices to find some $\mathcal{H}^{\prime}$ and $\mathcal{D}^{\prime}$ such that the following three conditions holds.
@@
    $\mathsf{Rad}_{n}(\mathcal{D}^{\prime},\mathcal{H}^{\prime}):=\frac{1}{n}\mathbb{E}_{\{x^{j}\}_{j=1}^{n}\overset{\mathrm{iid}}{\sim}\mathcal{D}^{\prime},\{\sigma^{j}\}_{j=1}^{n}\overset{\mathrm{iid}}{\sim}\{\pm 1\}}\left[\max_{h^{\prime}\in\mathcal{H}^{\prime}}\sum_{j=1}^{n}\sigma^{j}h^{\prime}(x^{j})\right]\leq C_{n};$ (117)

### v3

@@ E.2 Proof of Lemma 4
$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}=\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
=$\qquad\overset{\mathrm{(i)}}{=}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]\right]$
$\qquad\overset{\mathrm{(ii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
$\qquad=\left(\sum_{i=1}^{k}(n_{i}+m_{i})\right)\widetilde{\mathsf{Rad}}_{\{n_{i}+m_{i}\}_{i=1}^{k}}$
$\qquad\overset{\mathrm{(iii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
$\qquad\qquad\qquad\qquad\qquad+\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=n_{i}+1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
$\qquad=\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$
Here, (i) is valid due to the zero-mean property of $\{\sigma_{i}^{j}\}$, (ii) comes from Jensen’s inequality, and (iii) follows since $\max_{x}\big{(}f_{1}(x)+f_{2}(x)\big{)}\leq\max_{x}f_{1}(x)+\max_{x}f_{2}(x)$.
E.3 Proof of Lemma 5
Set $n=\sum_{i=1}^{k}n_{i}$. Let $\{X_{j}\}_{j=1}^{n}$ be $n$ i.i.d. multinomial random variables with parameter $\{w_{i}\}_{i=1}^{k}$, and take
$\widehat{n}_{i}=\sum_{j=1}^{n}\mathds{1}\{X_{j}=i\}$
for each $i\in[k]$. From (34) and Definition 2, it is easily seen that
$\mathsf{Rad}_{n}\big{(}D(w)\big{)}$ $=\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{n}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{n}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right]\right],$
@@ E.3 Proof of Lemma 5
with probability exceeding $1-1/(2k)$. Defining $\mathcal{E}$ to be the event that $\widehat{n}_{i}\geq n_{i}/6$ holds for all $i\in[k]$, we can invoke the union bound to see that
@@
$\mathsf{Rad}_{n}\big{(}\mathcal{D}(w)\big{)}$ $\geq\mathbb{P}(\mathcal{E})\cdot\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{n}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{n}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\mid\mathcal{E}\right]\right]\right]$
$\geq\frac{1}{2}\cdot\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{n}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widetilde{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\mid\mathcal{E}\right]\right]$
$=\frac{1}{2}\cdot\frac{\sum_{i=1}^{k}\widetilde{n}_{i}}{n}\widetilde{\mathsf{Rad}}_{\{\widetilde{n}_{i}\}_{i=1}^{k}}$
@@
thus concluding the proof.
E.4 Necessity of Assumption 1
In this subsection, we study whether Assumption 1 can be replaced by the following weaker assumption, the latter of which only assumes that the Rademacher complexity on each $\mathcal{D}_{i}$ is well-bounded.
@@ E.4 Necessity of Assumption 1
For each $n\geq 1$, there exists a quantity $\widetilde{C}_{n}>0$ (known to the learner) such that
$\widetilde{C}_{n}\geq\max_{1\leq i\leq k}\mathsf{Rad}_{n}(\mathcal{D}_{i}).$ (143)
@@
Let $w^{0}=[1/k,1/k,\ldots,1/k]^{\top}$. There exist a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that
$\mathsf{Rad}_{n}\big{(}\mathcal{D}(w^{0})\big{)}\geq\Omega\left(\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})\right)$ (144)
@@
Without loss of generality, consider the case where $\mathcal{Y}=\{0\}$ and $\ell(h,(x,y))=h(x)-y=h(x)$. We can then view $\mathcal{D}_{i}$ as a distribution over $\mathcal{X}_{i}$, as there is only one element in $\mathcal{Y}$.
Pick $k$ subsets of $\mathcal{X}$ as $\{\mathcal{X}_{i}\}_{i=1}^{k}$. For each $i\in[k]$, we choose the distribution $\mathcal{D}_{i}$ to be an arbitrary distribution supported on $\mathcal{X}_{i}$. In addition, we define $\mathcal{H}_{i}$ to be a set of hypothesis obeying $h(x)=0$ for all $x\notin\mathcal{X}_{i}$ for each $i\in[k]$. For a collection of hypothesis $\{h_{i}\}_{i=1}^{k}$ such that $h_{i}\in\mathcal{H}_{i}$, we define $\mathsf{joint}(\{h_{i}\}_{i=1}^{k})$ to be the hypothesis $h$ such that
$h(x)=\begin{cases}h_{i}(x)&\text{if }x\in\mathcal{X}_{i},~{}i\in[k];\\ 0&\text{if }x\notin\cup_{i}\mathcal{X}_{i}.\end{cases}$ (145)
The hypothesis set $\mathcal{H}$ is then constructed as
=$\mathcal{H}=\big{\{}\mathsf{joint}(\{h_{i}\}_{i=1}^{k})\mid h_{i}\in\mathcal{H}_{i},\forall i\in[k]\big{\}}.$ (146)
Recalling the definition of $\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$, we see that
$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$ $=\mathop{\mathbb{E}}\limits_{\{x_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h(x_{i}^{j})\right]\right]$
$=\mathop{\mathbb{E}}\limits_{\{x_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}\max_{h_{i}\in\mathcal{H}_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h_{i}(x_{i}^{j})\right]\right]$ (147)
@@
$=\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}n_{i}\mathsf{Rad}_{n_{i}}(\mathcal{D}_{i}),$
where (147) results from the definition of $\mathcal{H}$. By taking $n_{i}=\frac{n}{k}$ for all $i\in[k]$ and applying Lemma 5, we reach
$\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})=\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}\leq 72\mathsf{Rad}_{n}\big{(}\mathcal{D}(w^{0})\big{)}$ (148)
as claimed. ∎
By virtue of Lemma 21, if we set $\widetilde{C}_{n}=\widetilde{\Theta}(\sqrt{d/n})$ in Assumption 2, then the best possible upper bound on $\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))$ is $\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))=\widetilde{\Theta}(\sqrt{dk/n})$, which implies that more samples are needed to learn the mixed distribution $\mathcal{D}(w^{0})$ than learning each individual distribution. Moreover, under the construction in Lemma 21, if we further assume that $\min_{h_{i}\in\mathcal{H}_{i}}\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]=1/2$ for all $i\in[k]$, then to find $h$ such that
$\max_{1\leq i\leq k}\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{i}}[h(x)]\leq\frac{1}{2}+\varepsilon,$ (149)
we need to find, for each $i\in[k]$, a hypothesis $h_{i}\in\mathcal{H}_{i}$ such that
$\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]\leq\frac{1}{2}+\varepsilon.$ (150)
Following this intuition, we can construct a counter example under Assumption 2, with a formal theorem stated as follows.
@@
There exist a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that Assumption 2 holds with $\widetilde{C}_{n}=\widetilde{O}\left(\sqrt{\frac{d}{n}}\right)$, and it takes at least $\widetilde{\Omega}\left(\frac{dk}{\varepsilon^{2}}\right)$ samples to find some $h\in\mathcal{H}$ obeying
$\max_{i\in[k]}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\leq\min_{h^{\prime}\in\mathcal{H}}\max_{i\in[k]}L(h^{\prime},\mathrm{e}^{\mathsf{basis}}_{i})+\varepsilon.$
@@
With the construction in Lemma 21, it suffices to find some $\mathcal{H}^{\prime}$ and $\mathcal{D}^{\prime}$ such that the following three conditions hold:
@@
    The following inequality holds:
    $\mathsf{Rad}_{n}(\mathcal{D}^{\prime},\mathcal{H}^{\prime})\coloneqq\frac{1}{n}\mathop{\mathbb{E}}_{\{x^{j}\}_{j=1}^{n}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{D}^{\prime},\,\{\sigma^{j}\}_{j=1}^{n}\overset{\mathrm{i.i.d.}}{\sim}\{\pm 1\}}\left[\max_{h^{\prime}\in\mathcal{H}^{\prime}}\sum_{j=1}^{n}\sigma^{j}h^{\prime}(x^{j})\right]\leq\widetilde{C}_{n};$ (151)